Think about all the four - digit numbers you could make with the digits 1-4. What is the probabilty that the first digit will be a 3? Please help. Explain fully. Make me understand so I can explain to others.
Answer by mananth(16949)

1,2,3,4
The sample space is the set of all possible outcomes

S={1234, 1243, 1324, 1342, 1423, 1432, 2134, 2143, 2314, 2341, 2413, 2431, 3124
3142, 3214, 3241, 3412, 3421, 4123, 4132, 4213, 4231, 4312,4321,}
24 numbers can be formed.
number of sample points = n(S)=24
We should get a number starting from 3. How many are there?
Number of favourable outcomes Or event of a number starting from 3
denote by set A
A={3124, 3142, 3214, 3241, 3412, 3421,}
n(A)=6
probability of the number starting from 3 =( number of favorable outcomes/number of possible outcomes)
n(P) = n(A)/n(S)
=6/24
=1/4